OpenStudy (anonymous):

The ordered pair below is a solution to which equation? (2, 12) A. y = 6x B. C. y = 2x D.

OpenStudy (anonymous):

Confused. Please re post.

OpenStudy (anonymous):

The ordered pair below is a solution to which equation? (2, 12) A. y = 6x B. y=1/2x C. y = 2x D. y=1/6x

OpenStudy (anonymous):

Defiantly the first one.